OpenStudy (anonymous):

what is 3/6 divided by 5/9?

OpenStudy (anonymous):

For a division of 2 fractions a / b and c / d, (a / b) / (c / d) = (a / b)(d / c) = ad / (bc)

OpenStudy (anonymous):

Using what TURITW says:\[\frac{ 3 }{ 6 } \div \frac{ 5 }{ 9 }\]Flip and Multiply!\[\frac{ 3 }{ 6 } \times \frac{ 9 }{ 5 }\]Cancel out and solve! Can you do the rest?

OpenStudy (anonymous):

We got 27/30th or 10 1/30th? Is this correct?

OpenStudy (anonymous):

It's correct but you need to simplify it further

OpenStudy (anonymous):

27 / 30 is correct, not 10 1 / 30

OpenStudy (anonymous):

So what would the answer be?

OpenStudy (anonymous):

27 = 3^3 30 = 2 x 3 x 5 27 / 30 = 3^3 / (2(3)(5)) = 3^2 / (2(5)) = 9 / 10 Or, you could just type in 27 / 30 in a calculator

OpenStudy (anonymous):

9/10 is your answer

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Anytime. Let me know if your need more help!

OpenStudy (anonymous):